A wife and husband, Linda and Mark, have ages that total 80 years. Two years ago, 2 times Mark's age was 2 years more than 3 times Linda's age. How old is each now?
Answer by Paul(988)

Let Linda's age be x
Let Mark's age be y
x+y=80 (solve for y)
y=80-x (subitution)

Two years ago Mark's age = y-2
Two years ago Linda's age = x-2

Equation:
3(x-2)+2=2(y-2)
3x-6+2=2y-4
Subsitue for y
3x-4=2(80-x)-4
3x-4=160-2x-4
5x=160
x=32

y=80-32
y=48

Hence, Linda is 32 years old and Mark is 48 years old.
PAul.